Continuously Compounded Interest:
How much would $900 invested at 6% interest compounded continuously be worth after 4 years?

Answer by ankor@dixie-net.com(12692)

How much would $900 invested at 6% interest compounded continuously be worth after 4 years?
:
The continuous interest formula: A = P * e^rt
Where
P = amt invested
r = interest rate (in decimal form)
t = time in yrs
A = resulting amt
:
A = 900 * e^(.06*4)
A = 900 * e^.24
Find e^.24 with a calc
A = 900 * 1.27125
A = $1144.12